OpenStudy (anonymous):

Factor this expression completely, then place the factors in the proper location on the grid. 1/8x^3-1/27y^3

OpenStudy (anonymous):

\[(a-b)^3=(a-b)(a^2+ab+b^2)\] Find your a and b and use the formula.

OpenStudy (anonymous):

can you answer ?

OpenStudy (anonymous):

i give medal

OpenStudy (anonymous):

can you tell what a and b is? from your equation?

OpenStudy (anonymous):

\[a^3-b^3=(a-b)(a^2+ab+b^2)\]

OpenStudy (anonymous):

look on screen

OpenStudy (anonymous):

i understand the problem, but i'm helping you do it, not doing it for you friend.

OpenStudy (anonymous):

im stuck + this was last question of the day

OpenStudy (anonymous):

solve please

OpenStudy (anonymous):

its not hard, i can help you... can you tell me what a^3 is?

OpenStudy (anonymous):

a squared 3

OpenStudy (anonymous):

from the formula above, you need to find a and b using cube roots

OpenStudy (anonymous):

i hate fractions doing in factor

OpenStudy (anonymous):

\[\frac{ 1 }{ 8 }x^3-\frac{ 1 }{ 27 }y^3\] this is the formula for factoring: \[a^3-b^3=(a-b)(a^2+ab+b^2)\] You need to find the a and the b. what times itself 3 times gives 8 and what times itself gives 27?

OpenStudy (anonymous):

then pretend they're not fractions, just use the denominator for now pretend its 8x^3-27y^3

OpenStudy (anonymous):

what is the cube root of 8 and what is the cube root of 27?

OpenStudy (anonymous):

The cube root of 8 is 2

OpenStudy (anonymous):

good

OpenStudy (anonymous):

for 27 is 3

OpenStudy (anonymous):

good. so use the formula with a=2x and b=3y \[a^3-b^3=(a-b)(a^2-ab-b^2)\] \[8x^3-27y^3=(2x-3y)((2x)^2+(2x)(3y)+(3y)^2)\] The only difference with your problem is the numbers will be denominators with 1 as the numerator.

OpenStudy (anonymous):

i have to put 1/6x in grid with they answer

OpenStudy (anonymous):

yes the answer will have a 1/6 in it somewhere

OpenStudy (anonymous):

actually it should be 1/6xy

OpenStudy (anonymous):

1/216 (3x-2y)(9x^2+6xy+4y^2) that's what comes up to me

OpenStudy (anonymous):

\[8x^3−27y^3=(2x−3y)((2x)^2+(2x)(3y)+(3y)^2)\] \[(2x−3y)((2x)^2+(2x)(3y)+(3y)^2)\] is the part that needs calculating. The first two terms are done, now what is the first term of the second part of the parenthesis (2x)^2?

OpenStudy (anonymous):

4x^2?

OpenStudy (anonymous):

yes... so what's the rest?

OpenStudy (anonymous):

(4x^2 + 6xy + 9y^2) is the last part... therefore \[8x^3-27y^2=(2x-3y)(4x^2+6xy+9x^2)\] the only difference is now all the numbers are under the fraction with a 1 on top.

OpenStudy (anonymous):

8^3-6xy-27Y^3=9y^2

OpenStudy (anonymous):

?

OpenStudy (anonymous):

\[\frac{ 1 }{ 8}x^3-\frac{ 1 }{ 27}y^3=(\frac{ 1 }{ 2}x-\frac{ 1 }{ 3}y)(\frac{ 1 }{ 4}x^2+\frac{ 1 }{ 6}xy+\frac{ 1 }{ 9}y^2)\]

OpenStudy (anonymous):

thats the answer?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

their is no 1/8 in box look

OpenStudy (anonymous):

its the problem and the answer together

OpenStudy (anonymous):

i need put answer on grid can you type your answer so it will fit the gird look on screen

OpenStudy (anonymous):

ok the right of the equal sign is the answer, the left of the equal sign is the orginal problem

OpenStudy (anonymous):

ok got it :)

OpenStudy (anonymous):

are you teacher? in past or right now :) your good explaining

OpenStudy (anonymous):

yes i am just here to help people

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

its just a formula you plug the numbers in claylordmath, you're making it way harder than it is. i know it looks scary but its not really that hard.